Circular No. 8565

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) CBAT@CFA.HARVARD.EDU (science)

        URL http://cfa-www.harvard.edu/iau/cbat.html ISSN 0081-0304

Phone 617-495-7440/7244/7444 (for emergency use only)

SUPERNOVA 2005cs IN M51 W. Li and A. V. Filippenko, University of California, Berkeley; and S.

D. Van Dyk, Spitzer Science Center, California Institute of Technology, report: “New images from the Advanced Camera for Surveys (ACS) onboard the Hubble Space Telescope, obtained as part of GO-10182 (PI: Filippenko) on July 11.50 UT, confirm our previous identification (IAUC 8556) of the progenitor of the type-II SN 2005cs. When the new HRC/F250W images are geometrically transformed to the pre-supernova WFC/F814W images, the SN 2005cs position is consistent with our identified progenitor to within 0.1 ACS pixel (0′′.005).”

SUPERNOVA 2005cu IN NGC 6754 D. C. Leonard, California Institute of Technology, reports that CCD

spectra (range 330–900 nm) of SN 2005cu (IAUC 8562), obtained on July 12 UT with the Keck I 10-m telescope (+ LRISp), shows it to be a young type-II supernova. The spectrum consists of a blue continuum and P-Cyg profiles of the Balmer and He lines. Adopting the NED recession velocity of 3257 km/s for the host galaxy (from Mathewson et al. 1992, Ap.J. Suppl. 81, 413), the expansion velocity derived from the minimum of the Hβ line is ∼ 9000 km/s. Zero-velocity interstellar Na I D absorption with an equivalent width of ∼ 0.07 nm is detected in the spectrum of SN 2005cu, indicating gas along the line-of-sight in our galaxy and thus suggesting some reddening by dust. Schlegel et al. (1998, Ap.J. 500, 525) estimate E(B − V) = 0.07 mag of Galactic reddening along the line-of-sight to NGC 6754. Interstellar Na I D absorption is detected also at the redshift of the host galaxy, with an equivalent width of ∼ 0.13 nm, suggesting some host-galaxy extinction.

COMETS C/2001 Q9 AND C/2001 S3 (SOHO) Additional Kreutz sungrazing comets found on SOHO website images

(cf. IAUC 8563), both objects appearing stellar:

Comet 2001 UT α_{2000} δ_{2000} Inst. F MPEC
C/2001 Q9 Aug. 16.221 9^{h}24.^{m}3 +12^{o}32′ C3 RK 2005-N60
C/2001 S3 Sept. 28.763 12 03.3 − 2 57 C3 RK 2005-N60

X-RAY FLASH IN M51 Corrigendum. On IAUC 8564, line 20, for 4800 s read 48000 s

2005 July 12  \begin{matrix}c \\ ◯ \end{matrix} Copyright 2005 CBAT  Daniel W. E. Green